if carolines score is 5 less than twice Erins score it is also 40 points graiter than erins score.what score did the two girls recive

Answer by mananth(16946)

if carolines score is 5 less than twice Erins score it is also 40 points graiter than erins score.what score did the two girls recive
..
Erin's score =e
Carolina's score = 2e-5
...
2e-5=e+40
+5 +5
2e=e+45
-e -e
e=45 Erin's score.
carolina's score = 2e-5 = 2*45-5=85
...